# ACES: Automatic Cohort Extraction System for Event-Stream Datasets

### v2

@@ Abstract
Reproducibility remains a significant challenge in machine learning (ML) for healthcare. Datasets, model pipelines, and even task/cohort definitions are often private in this field, leading to a significant barrier in sharing, iterating, and understanding ML results on electronic health record (EHR) datasets. This paper addresses a significant part of this problem by introducing the Automatic Cohort Extraction System (ACES) for event-stream data. This library is designed to simultaneously simplify the development of task/cohorts for ML in healthcare and also enable the reproduction of these cohorts, both at an exact level for single datasets and at a conceptual level across datasets. To accomplish this, ACES provides (1) a highly intuitive and expressive configuration language for defining both dataset-specific concepts and dataset-agnostic inclusion/exclusion criteria, and (2) a pipeline to automatically extract patient records that meet these defined criteria from real-world data. ACES can be automatically applied to any dataset in either the Medical Event Data Standard (MEDS) or EventStreamGPT (ESGPT) formats, or to any dataset in which the necessary task-specific predicates can be extracted in an event-stream form. ACES has the potential to significantly lower the barrier to entry for defining ML tasks that learn representations, redefine the way researchers interact with EHR datasets, and significantly improve the state of reproducibility for ML studies in this modality. ACES is available at: https://github.com/justin13601/aces. A short video demonstration of ACES is available at: https://youtu.be/i_hCaHDydqA.
@@ 1 Introduction
Machine learning (ML) for healthcare suffers from a severe and systemic reproducibility crisis (McDermott et al., 2021b). This challenge is further exacerbated by the need to maintain private and secure datasets, but even with public datasets, ML pipelines are not reliably reproducible from published papers alone. For instance, in numerous attempts to reproduce ML for healthcare studies using the MIMIC-III dataset (Johnson et al., 2016), Johnson et al. found that more than half the time, the cohorts described in the studies could not be reliably reconstructed (Johnson et al., 2017).
@@
This burden in reproducing even the basic task and problem definitions in ML for healthcare studies is profoundly detrimental. Beyond the obvious concerns it raises around the robustness of reported results and their readiness for deployment, our inability to reliably define shared, canonical, reproducible task definitions limits our capacity to perform meaningful model comparisons during methodological development. This is particularly notable in settings where not all researchers have mutual access to all datasets, as is common in healthcare. Given the critical role that open benchmarks play in the advancement of ML methods (Zhang & Hardt, 2024; Salaudeen & Hardt, 2024; Shirali et al., 2023), this deficit directly translates to a significant barrier in our ability, as a research community, to effectively experiment, iterate, and develop new ML methodologies in the healthcare space.
@@
Given the clear import of this problem, the research community has naturally explored a number of prospective solutions. These can be largely categorized into two areas: (1) leveraging existing common data models (CDMs) to define reproducible task cohorts only for datasets within these schemas, and (2) defining static benchmarking tasks on individual public datasets. Both of these areas have generated numerous high-impact works. For example, in the area of CDM-driven tools, systems such as the ATLAS tool (OHDSI et al., 2024) for OHDSI’s OMOP CDM (OHDSI, 2021) and i2b2’s query tool for the i2b2 CDM (Murphy et al., 2010), as well as institution-specific tools such as the Stanford Research Repository (STARR) OMOP system and the NIHR Infections in Oxfordshire Database (IORD) cohort discovery platform have all been used to drive numerous new lines of inquiry. Unfortunately, these tools are also extremely limited in that they can only apply to the specific CDM or institutional data warehouse for which they have been defined. Further, because many of these CDMs have had limited penetrance into healthcare’s high-capacity, deep learning ecosystems, they are particularly ill-posed for task and cohort extraction within the healthcare deep learning communities. Conversely, public static benchmarks (Harutyunyan et al., 2019; McDermott et al., 2021a; van de Water et al., 2024) over datasets such as MIMIC-IV (Johnson et al., 2023) or eICU (Pollard et al., 2018) have also been extremely impactful for the ML community. However, they are all tied to only a single or small set of datasets and tasks. Given the highly dynamic nature of clinical data and healthcare requirements, this is insufficient for the benchmarking and reproducibility needs faced by the ML for healthcare community.
@@ 2. The Need for Flexibility
A single, static benchmark cannot encompass the variety of clinical tasks relevant to clinicians and informaticians. As existing tools (with only limited interfaces to define queries using per-set vocabularies) may struggle to generalize to new clinical tasks, ideal solutions must be sufficiently flexible to accommodate a myriad of new task definitions, criteria formats, and disease or deployment areas.
3. The Need for Accessibility, Usability, & Applicability in Deep Learning Workflows
While many existing tools present with no-code interfaces (e.g., web platforms to build queries) that are essential for less technically-literate audiences, integrating such tools with deep learning workflows can prove challenging. Deep learning systems are often run in a semi-programmatic manner on siloed, private computational clusters where researchers have minimal control. Hence, existing tools can cause significant hindrance. Instead, successful software must be able to provide a Python and command-line interface (CLI) that offer significant ease of use to deep learning researchers, alongside shareable and readable configuration files that specify task definitions in a manner that can be readily ported across datasets and environments.
@@ 3. The Need for Accessibility, Usability, & Applicability in Deep Learning Workflows
In this work, we solve these problems with the Automatic Cohort Extraction System for Event-Stream Datasets (ACES). ACES offers a simple, expressive, and shareable configuration language for task definitions, as well as a reliable command-line interface and Python-accessible library for extracting labeled task dataframes (Figure 1).
Task definitions in ACES are naturally separated into simple dataset-specific event predicates and dataset-agnostic inclusion/exclusion criteria, thereby permitting the same task to be used in a conceptually identical manner across diverse datasets. By requiring users to specify predicates to realize their ML tasks on their specific datasets, ACES allows users to produce precise, locally specific, validatable cohorts that harmonize only the data elements needed for their task, regardless of how their input dataset is aligned or misaligned with existing ontologies or CDMs. Further, for datasets that are fully harmonized (e.g., through OHDSI vocabularies), ACES predicate definitions can be re-used across datasets without any loss of utility. In this way, ACES accommodates diverse datasets at various levels of data harmonization in a flexible, transparent manner. Overall, this approach not only enhances reproducibility but also facilitates community collaboration on task definitions, inclusion/exclusion criteria, and evaluation metrics for specific clinical use cases.
In contrast to prior task-definition systems such as ATLAS, ACES makes minimal assumptions about the input data structure or source vocabularies. In particular, ACES can be run on any dataset, provided the necessary task-specific predicates can be pre-extracted in an ”event-stream” format111e.g., in a dataframe with columns for a subject ID, timestamp, and numerical predicate values., and can further be run from raw data directly for any dataset in the relatively low-level and flexible Medical Event Data Standard (MEDS) (Arnrich et al., 2024) or EventStreamGPT (ESGPT) (McDermott et al., 2023) formats. To demonstrate the utility and flexibility of ACES, we also release a collection of task definitions (Section 4) based on prior ML for healthcare works at both the dataset-agnostic criteria level and with dataset-specific predicates for the widely-used MIMIC-IV dataset (Johnson et al., 2023).
@@
    Through these advancements, ACES is poised to significantly advance the state of reproducibility, interoperability, and effective development of ML methods for healthcare.
In the rest of this work, we will do the following. First, in Section 2, we will go through ACES in more depth, beginning by illustrating the key concepts of ACES using a running example with ACES CLI. Next, in Section 3, we will discuss the ACES design principles and briefly overview its core algorithm. Subsequently, in Section 4, we will demonstrate the use of ACES on diverse problem areas over real-world data, before ultimately discussing the limitations and future roadmap of ACES in Section 5, and offering concluding thoughts in Section 6.
Figure 1: ML task cohort extraction process (A) with and (B) without ACES. Predicates are dataset-specific concepts that are needed to conceptually capture a machine learning task. Windows are temporal segments on a patient’s health record and are dataset-agnostic, as they are defined relative to the predicates. This distinction allows researchers to share the more complex task logic that is independent of datasets, enabling conceptual reproducibility for ML tasks in healthcare.
@@ 2 Automatic Cohort Extraction System for Event-Stream Datasets (ACES)
In this section, we introduce ACES, a novel automated task and cohort definition and extraction system that fills the key gaps in interoperability, flexibility, and accessibility left by the existing tools outlined in Section 1.
To use ACES and extract a cohort for downstream ML tasks, a user only needs to do the following simple steps:
1. Install ACES:
@@
A fully functional version of ACES is pushed to PyPI, and any user can easily install it by simply running pip install es-aces. All dependencies are automatically set up with no further actions needed.
2. Define Dataset:
@@
A dataset in a permitted format, such as MEDS, ESGPT, or as direct predicates, is required. More information on the data formats is available in Section 2.1.
3. Define Task:
@@
A task configuration file is required to define the task that the user wishes to extract. This configuration language is simple, clear, yet flexible, permitting users to rapidly share and iterate over task definitions for their clinical settings. Configuration specification is given in Section 2.2 and the Appendix.
4. Run the ACES CLI:
ACES can be directly run from the command-line using:
1\$ aces-cli cohort_name="\$TASK" data.path="\$DATA_PATH"
@@ 4. Run the ACES CLI:
Additional details about the possible command-line arguments are detailed in Section 2.3. ACES can also be used as a direct Python import, as detailed in Section 2.5.
5. Get Outputs:
@@
The resulting output from ACES is a single unified dataframe with all valid patient instances extracted according to task specifications. Users can subsequently leverage the returned columns with original patient identifiers and health record timestamps for downstream ML tasks and benchmark creation. ACES also returns additional columns for the task label, as well as summaries of each of the predicates over windows of the patient record. Further information is available in Section 2.4 and in the example output displayed in Figure 3.
@@
Critically, after only these five simple steps, a user can immediately, reproducibly extract a full cohort from their source dataset that matches their task definition, and begin using this task for downstream representation learning. Due to space constraints, full technical details on all aspects of ACES and on the precise details of the recursive algorithm used to extract task cohorts given a configuration file are limited to the Appendix and the ACES public documentation, available here: https://eventstreamaces.readthedocs.io/en/latest/technical.html. Please also refer to the PDF compilation of the technical documentation in the Supplementary Material (Pages 41-51).
@@ 2.1 Dataset Configurations
ACES is extremely flexible and can handle different data formats, including data.standard=meds (Arnrich et al., 2024), data.standard=esgpt (McDermott et al., 2023), or data.standard=direct, where event-stream predicate values are pre-extracted by the user from any given dataset schema. Other CDMs are interchangeable with these formats, such as OHDSI OMOP through the MEDS OMOP ETL [^2] https://github.com/Medical-Event-Data-Standard/meds_etl/tree/main.
The use of direct predicates for extraction from datasets in a format that ACES does not natively support still significantly reduces the burden on users. Simply creating predicates is much less cumbersome than either fully converting the dataset to a CDM in order to use existing tools like ATLAS or i2b2, or performing the entire task extraction from scratch by writing in-house dataframe querying code. This demonstrates the significant improvement in utility that ACES brings across diverse data schemas compared to existing tools.
@@ 2.2 Task Configurations
In ACES, tasks are specified through configuration files that define a collection of dataset-specific event predicates, which are simple functions evaluated on individual events within a structured event-stream dataset. Additionally, task criteria are defined in a dataset-agnostic manner through a collection of interrelated windows, which specify regions of a patient’s record and are constrained by certain relationships. Please see Figure 2 for an example of a task configuration.
Figure 2: Example of a configuration file for the binary prediction of in-hospital mortality 48 hours after admission. References to predicates and windows are italicized and bolded, respectively. (A) Dataset-specific task predicates. These concepts are needed to conceptually capture this task and are used as constraints and boundaries for windows of the patient record. For instance, in this example, the value of “ADMISSION” denotes a hospital admission event in the external source dataset. (B) A window of the task specifying the task inputs for downstream models. Suppose we’d like to use all historic patient data up to and including 24 hours past the admission. An arbitrary criterion requiring more than 5 records can then be placed on this window to ensure that the extracted cohort contains sufficient input data. (C) The trigger events for the task, which are hospital admissions as we’d like to make a mortality prediction for each admission. (D) A window of the task specifying a gap in the patient timeline. Suppose we’d like to set a minimum length of admission for our cohort (e.g., 48 hours). A temporal constraint (minimum window duration) of 48 hours could then be set to represent this requirement. (E) A window of the task specifying the task target, which is set from the end of (D) to the immediately subsequent “discharge” or “death” predicate. This creates our binary classes for the task (i.e., $discharge=0$, $death=1$). All windows are interrelated on the patient timeline, as shown by how each window references another in the configuration file.
2.3 Command-Line Interface Arguments

### v3

@@ Abstract
Reproducibility remains a significant challenge in machine learning (ML) for healthcare. Datasets, model pipelines, and even task or cohort definitions are often private in this field, leading to a significant barrier in sharing, iterating, and understanding ML results on electronic health record (EHR) datasets. We address a significant part of this problem by introducing the Automatic Cohort Extraction System (ACES) for event-stream data. This library is designed to simultaneously simplify the development of tasks and cohorts for ML in healthcare and also enable their reproduction, both at an exact level for single datasets and at a conceptual level across datasets. To accomplish this, ACES provides: (1) a highly intuitive and expressive domain-specific configuration language for defining both dataset-specific concepts and dataset-agnostic inclusion or exclusion criteria, and (2) a pipeline to automatically extract patient records that meet these defined criteria from real-world data. ACES can be automatically applied to any dataset in either the Medical Event Data Standard (MEDS) or Event Stream GPT (ESGPT) formats, or to any dataset in which the necessary task-specific predicates can be extracted in an event-stream form. ACES has the potential to significantly lower the barrier to entry for defining ML tasks in representation learning, redefine the way researchers interact with EHR datasets, and significantly improve the state of reproducibility for ML studies using this modality. ACES is available at: https://github.com/justin13601/aces.
@@ 1 Introduction
Machine learning (ML) for healthcare suffers from a severe and systemic reproducibility crisis (McDermott et al., 2021b). This challenge is further exacerbated by the need to maintain private and secure datasets, but even with public datasets, ML pipelines are not reliably reproducible from published papers alone. For instance, in numerous attempts to reproduce ML for healthcare studies using the MIMIC-III dataset (Johnson et al., 2016), Johnson et al. found that more than half the time, the cohorts described in the studies could not be reliably reconstructed. Specifically, experiments led to many discrepancies of up to 25% in cohort sizes, with one study reaching as high as 11,767 patients (Johnson et al., 2017). This is primarily due to sparse descriptions of cohorts in study methods with essential reproducibility details often omitted, along with the absence of openly available code.
@@
This burden in reproducing even the basic task and problem definitions in ML for healthcare studies is profoundly detrimental (McDermott, 2025). Beyond the obvious concerns it raises around the robustness of reported results and their readiness for deployment, our inability to reliably define shared, canonical, and reproducible task definitions limits our capacity to perform meaningful model comparisons during methodological development. This is particularly notable in settings where not all researchers have mutual access to all datasets, as is common in healthcare. Given the critical role that open benchmarks play in the advancement of ML methods (Zhang & Hardt, 2024; Salaudeen & Hardt, 2024; Shirali et al., 2023), this deficit directly translates to a significant barrier in our ability, as a research community, to effectively experiment, iterate, and develop new ML methodologies in the healthcare space.
@@
Given the clear import of this problem, the research community has naturally explored a number of prospective solutions. These can largely be categorized into two areas: (1) leveraging existing common data models (CDMs) to define reproducible task cohorts only for datasets within these schemas, and (2) defining static benchmarking tasks on individual public datasets. Both of these areas have generated numerous high-impact works. For example, in the area of CDM-driven tools, systems such as the ATLAS tool (Gold et al., 2024) for OHDSI’s OMOP CDM (Reich et al., 2024) and i2b2’s PIC-SURE (Stedman et al., 2024) for the i2b2 CDM (Murphy et al., 2010), as well as various institution-specific tools, have all been used to drive numerous new lines of inquiry. Unfortunately, these tools are also extremely limited in that they can only be applied to the specific CDM or institutional data warehouse for which they have been defined. Further, because many of these CDMs have had limited penetrance into healthcare’s high-capacity, deep learning ecosystems, they are particularly ill-posed for task and cohort extraction within the healthcare deep learning communities. Conversely, public static benchmarks (Harutyunyan et al., 2019; McDermott et al., 2021a; van de Water et al., 2024) over datasets such as MIMIC-IV (Johnson et al., 2023) or eICU (Pollard et al., 2018) have also been extremely impactful. However, they are all tied to only a single or small set of datasets and tasks. Given the highly dynamic nature of clinical data and healthcare requirements, this is insufficient for the benchmarking and reproducibility needs faced by the ML for healthcare community.
@@ 2. The Need for Flexibility
A single, static benchmark cannot encompass the variety of clinical tasks relevant to clinicians and informaticians. As existing tools (with limited interfaces for defining queries using per-set vocabularies) may struggle to generalize to new clinical tasks, ideal solutions must be sufficiently flexible to accommodate a myriad of new task definitions, criteria formats, and disease or deployment areas.
3. The Need for Accessibility, Usability, and Applicability in Deep Learning Workflows
While many existing tools feature no-code interfaces (e.g., web platforms to build queries) that are essential for less technically-literate audiences, integrating such tools with deep learning workflows can prove challenging. Deep learning systems are often run in a semi-programmatic manner on siloed, private computational clusters where researchers have minimal control. Hence, existing tools can cause significant hindrance. Instead, successful software must be able to provide a Python and command-line interface (CLI) that offer significant ease of use to deep learning researchers, alongside shareable and readable configuration files that specify task definitions in a manner that can be readily ported across datasets and environments.
@@ 3. The Need for Accessibility, Usability, and Applicability in Deep Learning Workflows
In this work, we solve these problems with the Automatic Cohort Extraction System for Event-Stream Datasets (ACES). ACES offers a simple, expressive, and shareable configuration language for task and cohort definitions, as well as a reliable CLI and a straightforward Python library for extracting labeled task dataframes (Figure 2).
Task definitions in ACES are naturally separated into simple dataset-specific event predicates and dataset-agnostic inclusion or exclusion criteria, thereby permitting the same task to be used in a conceptually identical manner across diverse datasets. By requiring users to specify predicates to realize their ML tasks on their specific datasets, ACES allows users to produce precise, locally-specific, verifiable cohorts that harmonize only the data elements needed for their task, regardless of how their input dataset is aligned or misaligned with existing ontologies or CDMs. Further, for datasets that are fully harmonized (e.g., through OHDSI vocabularies), ACES predicate definitions can be re-used across datasets without any loss of utility. In this way, ACES accommodates diverse datasets at various levels of data harmonization in a flexible, transparent manner. Overall, this approach not only enhances reproducibility but also facilitates community collaboration on task definitions, inclusion or exclusion criteria, and evaluation metrics for specific clinical use cases.
In contrast to prior task definition systems such as ATLAS, ACES makes minimal assumptions about the input data structure or source vocabularies. In particular, ACES can be run on any dataset, provided the necessary task-specific predicates can be pre-extracted in an “event-stream” format (Figure 1). It can further be run from raw data directly for any dataset in the relatively low-level and flexible Medical Event Data Standard (MEDS) (Arnrich et al., 2024) or Event Stream GPT (ESGPT) (McDermott et al., 2023) formats in approximately five lines of template code, offering high efficiency.
Figure 1: Workflow for extracting cohorts using ACES. The pipeline shows the expected format for ACES-supported event-stream datasets and outcome cohorts. The transformation of raw data into the event-stream format is intentionally designed to be straightforward — primarily merging relational database tables — minimizing data loss risks associated with other CDMs like OMOP.
Further, we align ACES with the concept of frictionless reproducibility for shared tasks proposed by David Donoho (Donoho, 2024), especially for the clinical domain. This addresses the “Bring-Your-Own-Data Challenge”, where much research relies on private patient outcomes data, often accessible only to a few credentialed researchers under strict usage agreements. Even when benchmarking platforms and shared code exist, the inability to share data directly remains a significant barrier and often stifles progress. ACES seeks to overcome these challenges by offering a domain-specific language (DSL) and novel infrastructure to ensure reproducibility without necessitating data sharing. Instead of relying on public datasets or reconfiguring code for diverse environments, ACES enables researchers to distribute task definitions through configuration files. These files provide a standardized way to conceptually reproduce cohorts on private datasets or exactly reproduce them on public datasets.
@@
    ACES introduces a novel DSL that leverages event-bounded query aggregations, enabling more expressive and efficient task definitions that better capture clinical timelines.
In the rest of this work, we first explore ACES in depth in Section 2, beginning with an illustration of its key concepts before briefly overviewing its core recursive algorithm. We then present a running example to illustrate how configuration files capture task logic. Next, in Section 3, we demonstrate the use of ACES across diverse problem areas using real-world data, releasing a collection of task definitions based on prior ML for healthcare works — both at the dataset-agnostic criteria level and with dataset-specific predicates for the widely-used MIMIC-IV dataset (Johnson et al., 2023). Finally, we discuss the limitations and future roadmap of ACES in Section 4, and offer concluding thoughts in Section 5.
Figure 2: ML task cohort extraction process (A) with and (B) without ACES. Predicates are dataset-specific concepts that are needed to conceptually capture a ML task. Windows are temporal segments on a patient’s health record and are dataset-agnostic, as they are defined relative to the predicates. This distinction allows researchers to easily share the more complex task logic which is independent of datasets, facilitating conceptual reproducibility for ML tasks in healthcare.
@@ 2 Automatic Cohort Extraction System for Event-Stream Datasets (ACES)
In this section, we introduce ACES, a novel automatic task and cohort extraction system that fills the key gaps in interoperability, flexibility, and accessibility left by the existing tools outlined in Section 1. To use ACES and extract a cohort for downstream ML tasks, a user only needs to do the following simple steps:
@@
1. Install ACES: A fully functional version of ACES is pushed to PyPI, and any user can easily install it by simply running pip install es-aces. All dependencies are automatically set up with no further actions needed.
@@
2. Define Dataset: A dataset in a permitted format, such as MEDS, ESGPT, or as direct predicates, is required. More information on the data formats is available in Section 2.2.
@@
3. Define Task: A task configuration file is required to define the task that the user wishes to extract. This configuration language is simple, clear, yet flexible, permitting users to rapidly share and iterate over task definitions for their clinical settings. Configuration specification is given in Section 2.3.
@@
4. Run the ACES CLI: ACES can be directly run from the command line. Details about the possible command-line arguments are detailed in Section 2.4. ACES can also be used as a direct Python import, as mentioned in Section 2.5.
Sample command for basic cohort extraction using ACES CLI:
1\$ aces-cli cohort_name='\$TASK' data.path='\$DATA_PATH'
@@
5. Get Outputs: ACES outputs a single unified dataframe with all valid patient instances extracted according to task specifications. Users can subsequently leverage the returned columns with original patient identifiers and health record timestamps for downstream ML tasks. Further information is available in Section 2.6.
@@
Critically, after only these five simple steps, a user can immediately, reproducibly extract a full cohort from their source dataset that matches their task definition, and begin using this task for downstream representation learning.
2.1 Algorithm Design
ACES addresses the challenge of extracting meaningful windows of data from patient records by using a recursive approach grounded in a tree-structured configuration file. Each task is represented as a hierarchy of constraints, with nodes defining boundaries of windows of interest and edges specifying temporal or event-based relationships between these windows. The algorithm begins by identifying root anchor events in the dataset that correspond to the triggering criteria of the task. It then recursively evaluates subtrees of constraints, aggregating predicate counts over defined windows either through temporal aggregations (e.g., over fixed time intervals) or event-bound aggregations (e.g., over windows bounded by specific clinical events, such as admissions, diagnoses, etc.). Each step ensures that the criteria of the subtree are met, filtering out invalid realizations before proceeding to child nodes. This recursive process guarantees that the specified configuration can always be resolved into valid windows that meet the task’s constraints. The final output is a dataframe containing all valid patients, task-specific labels, and prediction timestamps, and optionally, window start and end times as well as aggregated predicate counts. This ensures systematic and deterministic extraction of datasets for ML tasks. It also maintains flexibility and leverages the simple, transparent, and highly expressive DSL of ACES (Figure 3).
Figure 3: Overview of the ACES recursive algorithm. Given a task tree generated from a configuration file, ACES first identifies possible roots of the tree (task triggers) based on the associated predicate. It then computes aggregations of predicate counts over time-based (i.e., windows with a time interval) or event-based (i.e., windows between specified events) periods to summarize predicates over the edges between the tree nodes. Finally, invalid branches are filtered out if their predicate counts do not meet the specified criteria. This process is recursed for all child nodes of the task tree.
@@ 2.2 Dataset Configurations
ACES is extremely flexible and can handle different input data formats, including MEDS (data.standard=meds), ESGPT (data.standard=esgpt), or direct predicates (data.standard=direct), where event-stream features are pre-extracted by the user from any given dataset schema. Other CDMs are interchangeable with these formats, such as OHDSI OMOP through the MEDS OMOP ETL [^1] https://github.com/Medical-Event-Data-Standard/meds_etl/tree/main, which transforms OMOP-compliant datasets into MEDS without data loss or scalability issues.
Using direct predicates to extract cohorts from formats that ACES does not natively support still significantly reduces the burden on users. Simply creating predicate features is much less cumbersome than either fully converting the dataset to a CDM in order to use existing tools like ATLAS or i2b2’s platform, or performing the entire task extraction from scratch by writing in-house dataframe querying code. Additionally, as ACES configuration files are shareable and easily portable to other datasets (by simply swapping out predicate definitions), we believe ACES will offer long-term efficiency benefits. This demonstrates the significant improvement in utility that ACES brings across diverse data schemas compared to existing tools.
@@ 2.3 Task Configurations
In ACES, tasks are specified through configuration files that define a collection of dataset-specific event predicates, which are simple functions evaluated on individual events within a structured event-stream dataset. Predicate definitions can be stored in a central “database” file specific to each dataset, such that previously defined features could be easily reused for a variety of downstream tasks without further effort. Community predicate contributions for public datasets also streamline collaborative efforts for reproducibility. Additionally, task criteria are defined in a dataset-agnostic manner through a collection of interrelated windows, which specify segments of a patient’s record and are constrained by certain relationships. Please see Figure 4 for an example of a task configuration.
2.4 Command-Line Interface